OpenStudy (kj4uts):

A chemist has a 55.0 mL of a 1.50 M solution of NaOH available. Yet he needs 1.00 M solution. How many mL of 1.00 M can he make if he uses all the 55.0 mL? Remember: C1 X V1 = C2 X V2 A. 82.5 mL B. 40.0 mL C. 12.1 mL D. 0.0825 mL Please explain. Thank you!

OpenStudy (kj4uts):

I guess it is set up like this: 1.50 M X V1 = 1.00 M X 55 mL?

OpenStudy (abhisar):

\(\sf1.50 \times 55 mL = 1.00 \times V_2\)

OpenStudy (abhisar):

Got it?

OpenStudy (kj4uts):

Ok is see so then the 1.50 X 55 = 82.5 = 1.00 X V2. How do I solve that part?

OpenStudy (abhisar):

\(\sf \Rightarrow V_2=\frac{82.5}{1}=82.5\)

OpenStudy (kj4uts):

Oh I see so it is just 82.5 then.

OpenStudy (abhisar):

Yeppers!

OpenStudy (kj4uts):

Just curious does the 82.5 come out as mL?

OpenStudy (abhisar):

Obviously. You too 55.0 mL so the other volume will also be in mL. Take it as 0.055 L and the volume will come in L

OpenStudy (abhisar):

Cool?

OpenStudy (kj4uts):

Gotcha, thanks for your time and help!